# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

### v1

@@ Abstract
We consider the problem of conformal prediction under covariate shift. Given labeled data from a source domain and unlabeled data from a covariate shifted target domain, we seek to construct prediction sets with valid marginal coverage in the target domain. Most existing methods require estimating the unknown likelihood ratio function, which can be prohibitive for high-dimensional data such as images. To address this challenge, we introduce the likelihood ratio regularized quantile regression (LR-QR) algorithm, which combines the pinball loss with a novel choice of regularization in order to construct a threshold function without directly estimating the unknown likelihood ratio. We show that the LR-QR method has coverage at the desired level in the target domain, up to a small error term that we can control. Our proofs draw on a novel analysis of coverage via stability bounds from learning theory. Our experiments demonstrate that the LR-QR algorithm outperforms existing methods on high-dimensional prediction tasks, including a regression task for the Communities and Crime dataset, and an image classification task from the WILDS repository.
@@ 1 Introduction
  • We demonstrate the effectiveness of the LR-QR algorithm on high-dimensional datasets under covariate shift, including the Communities and Crime dataset and the RxRx1 dataset from the WILDS repository. Here, we leverage our theoretical insights, by choosing the regularization parameter by a cross validation procedure guided by theory. An implementation of LR-QR can be accessed at the following link: https://github.com/shayankiyani98/LR-QR.
@@ 1.1 Related work
Here we only list prior work most closely related to our method; we provide more references in Appendix C. The early ideas of conformal prediction were developed in Saunders et al., (1999); Vovk et al., (1999). With the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing prediction sets (e.g., Papadopoulos et al., , 2002; Vovk et al., , 2005; Vovk, , 2013). Classical conformal prediction guarantees validity when the calibration and test data are drawn from the same distribution. In contrast, when there is distribution shift between the calibration and test data (e.g., Quiñonero-Candela et al., , 2009; Shimodaira, , 2000; Sugiyama and Kawanabe, , 2012; Ben-David et al., , 2010; Taori et al., , 2020), coverage may not hold. Covariate shift is a type of dataset shift that arises in many settings, e.g., when predicting disease risk for individuals whose features may evolve over time, while the outcome distribution conditioned on the features remains stable (Quiñonero-Candela et al., , 2009).
@@ 3 Algorithmic principles
Here we present some intuition behind our approach.
3.1 Change of measure
A key fact for us is the change-of-measure identity $\mathbb{E}_{1}[r(X)h(X)]=\mathbb{E}_{2}[h(X)]$ for all measurable functions $h:\mathcal{X}\to\mathbb{R}$. To see this, we simply write $\mathbb{E}_{1}[r(X)h(X)]=\mathbb{E}_{1}[d\mathbb{P}_{2,X}/d\mathbb{P}_{1,X}(X)\cdot h(X)]=\mathbb{E}_{2}[h(X)]$. Therefore, for a threshold function $q$,
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant q(X)]]$ $=\mathbb{E}_{1}[r(X)\mathbb{P}_{Y|X}[S(X,Y)\leqslant q(X)]]$
$=\mathbb{E}_{2}[\mathbb{P}_{Y|X}[S(X,Y)\leqslant q(X)]]=\mathbb{P}_{2}[S(X,Y)\leqslant q(X)].$
Thus a threshold function $q$ yields a conformal predictor with marginal $(1-\alpha)$-coverage in the test domain if and only if
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant q(X)]]\geqslant 1-\alpha.$
3.2 Warm-up: stationary conditions of pinball loss
As a warm-up, we begin with a simpler objective function, the pinball loss. Suppose that we knew the true likelihood ratio $r$. Then in order to obtain our threshold function, we could solve the optimization problem $\kappa^{*}=\arg\min_{\kappa\in\mathbb{R}}\mathbb{E}_{1}[\ell_{\alpha}(\kappa r(X),S)],$ and set our threshold function to equal $\zeta^{*}=\kappa^{*}r$. We now argue that the marginal $(1-\alpha)$-coverage of $\zeta^{*}$ in the test domain can be deduced from the first order condition with respect to $\kappa$. Assuming that the distribution of $S|X=x$ has a density for all $x\in\mathcal{X}$, and using the fact that $\frac{\partial}{\partial c}\ell_{\alpha}(c,s)=\mathbf{1}[s\leqslant c]-(1-\alpha)$ for all $c\neq s$, the derivative of the objective with respect to $\kappa$ is
$\frac{\partial}{\partial\kappa}\mathbb{E}_{1}[\ell_{\alpha}(\kappa r(X),S)]$ $=\mathbb{E}_{1}\left[\frac{\partial}{\partial\kappa}\ell_{\alpha}(\kappa r(X),S)\right]$
$=\mathbb{E}_{1}[r(X)(\mathbf{1}[S\leqslant\kappa r(X)]-(1-\alpha))].$
Setting this equal to zero, we deduce that the minimizer $\kappa^{*}$ satisfies $\mathbb{E}_{1}[r(X)(\mathbf{1}[S\leqslant\kappa^{*}r(X)]-(1-\alpha))]=0$. By the change-of-measure identity above, this implies the desired marginal coverage guarantee for $\zeta^{*}$.
The above problem can be recast as optimizing over the one-dimensional function class ~=$\tilde{\mathcal{H}}=\ \{\kappa r:\kappa\in\mathbb{R}\}$ spanned by $r$. So, if we knew $r$, we could ensure coverage by optimizing over a single scalar parameter, which would be convenient and also have good finite sample coverage. However, we do not know $r$. Therefore, we consider a general class $\mathcal{H}$ large enough to approximate $r$ in a suitable way, and solve an analogous optimization problem. Intriguingly, the first order conditions will ensure coverage.
Specifically, given a convex function class $\mathcal{H}$, we consider the optimization problem
$(\tilde{h}^{*},\kappa^{*})=\arg\min_{\tilde{h}\in\mathcal{H},\kappa\in\mathbb{R}}\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h},S)],$
and set our threshold function equal to $\kappa^{*}\tilde{h}^{*}$. To derive the first order condition with respect to $\tilde{h}$, assume for simplicity that the minimizer $\tilde{h}^{*}$ lies in the interior of $\mathcal{H}$. Given a valid perturbation direction [^1] Formally, an element of the tangent cone of $\mathcal{H}$ at $\tilde{h}^{*}$. $g$, the function $\tilde{h}^{*}+\varepsilon g$ is in $\mathcal{H}$ for sufficiently small $\varepsilon$. Since for $\kappa=\kappa^{*}$, $\tilde{h}^{*}$ minimizes the objective in $\tilde{h}$, the function $\varepsilon\mapsto\mathbb{E}_{1}[\ell_{\alpha}(\kappa^{*}(\tilde{h}^{*}(X)+\varepsilon g(X)),S)]$ has a minimum at $\varepsilon=0$. Setting the derivative of this function with respect to $\varepsilon$ equal to zero as in Gibbs et al., (2023), we obtain the first order condition
@@ 3.2 Warm-up: stationary conditions of pinball loss
$\frac{\partial}{\partial\varepsilon}\bigg{|}_{\varepsilon=0}\mathbb{E}_{1}[\ell_{\alpha}(\kappa^{*}(\tilde{h}^{*}(X)+\varepsilon g(X)),S)]$ $=\mathbb{E}_{1}[\kappa^{*}g(X)(\mathbf{1}[S\leqslant\kappa^{*}\tilde{h}^{*}(X)]-(1-\alpha))]$
$=0.$
If $\kappa^{*}\neq 0$ and $\mathbb{E}_{1}[g(X)]\neq 0$, then by rearranging and dividing by $\kappa^{*}\mathbb{E}_{1}[g(X)]$, this is equivalent to $\mathbb{E}_{1}\left[\frac{g(X)}{\mathbb{E}_{1}[g(X)]}\mathbf{1}[S\leqslant\kappa^{*}\tilde{h}^{*}(X)]\right]=1-\alpha$. If the likelihood ratio $r$ is a valid perturbation at $\tilde{h}^{*}$, we can plug in $g=r$ to find that the threshold function $\kappa^{*}\tilde{h}^{*}$ attains valid coverage in the test domain. Thus, we search for classes $\mathcal{H}$ such that $r$ is a valid perturbation at all interior points of $\mathcal{H}$.
3.3 Constrained quantile regression
We now design a suitable function class $\mathcal{H}$. Given $\varepsilon>0$, consider in this section $\mathcal{H}=B_{\varepsilon}(r)$, the ball of radius $\sqrt{\varepsilon}$ centered at $r$, consisting of all square-integrable functions $\tilde{h}$ satisfying $\mathbb{E}_{1}[(\tilde{h}(X)-r(X))^{2}]\leqslant\varepsilon$. Since the interior is an open set, all perturbations are valid. Thus, by our remarks above, if we minimize $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]$ for scalar $\kappa$ and $\tilde{h}$ restricted to this ball, we obtain a threshold function with valid coverage under the covariate shift defined by $r$.
3.4 Removing explicit dependence on the likelihood ratio
Next, notice that we can expand $\mathbb{E}_{1}[(\tilde{h}(X)-r(X))^{2}]$ into
$\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{1}[-2r(X)\tilde{h}(X)]+\mathbb{E}_{1}[r(X)^{2}]=\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)]+\mathbb{E}_{1}[r(X)^{2}],$
@@ 3.4 Removing explicit dependence on the likelihood ratio
where we used the change-of-measure identity $\mathbb{E}_{1}[r(X)\tilde{h}(X)]=\mathbb{E}_{2}[\tilde{h}(X)]$. Only the first two terms depend on the optimization variable $\tilde{h}$, so it is enough to consider these two terms. This is where we make our second key observation. The insight is that neither of the terms $\mathbb{E}_{1}[\tilde{h}(X)^{2}]$ or $\mathbb{E}_{2}[-2\tilde{h}(X)]$ explicitly involve $r$, and thus they can be estimated by sample averages over the source and target data, respectively. Thus, we can minimize $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]$ over $\kappa\in\mathbb{R}$ while keeping $\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)]$ bounded. The threshold $\kappa^{*}\tilde{h}^{*}$ will have valid coverage under the covariate shift $r$.
@@ 3.5 Algorithm: likelihood ratio regularized quantile regression
Further, we can equivalently replace the constraint with a regularization term. Given a regularization strength $\lambda\geqslant 0$, we can minimize the objective $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]+\lambda(\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)])$ over $\kappa\in\mathbb{R}$ and $\tilde{h}$ square integrable with respect to $\mathbb{P}_{1,X}$. By reparametrizing this optimization problem ($\tilde{h}\mapsto\beta h,\kappa\mapsto 1/\beta$), we may move the scalar parameter $\kappa$ from the first term to the second and third terms. This yields the population-level likelihood ratio regularized quantile regression (LR-QR) objective
$L(h,\beta)=\mathbb{E}_{1}[\ell_{\alpha}(h,S)]+\lambda\mathbb{E}_{1}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h],$ (2)
for some regularization strength $\lambda\geqslant 0$. Finally, since in practice, we cannot optimize over all square-integrable functions, we restrict to functions $h$ in some function class $\mathcal{H}$, leading to the scaled LR-QR problem
$(h^{*},\beta^{*})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}L(h,\beta).$ (LR-QR)
We set our threshold function $q$ in (1) to equal $h^{*}$, so that our conformal prediction sets equal $C^{*}(x)=\{y\in\mathcal{Y}:S(x,y)\leqslant h^{*}(x)\}$ for all $x\in\mathcal{X}$.
We solve an empirical version of this objective. We will use our labeled source data $\{(X_{i},Y_{i}):i\in[n_{1}]\}$ to estimate $\mathbb{E}_{1}[\ell_{\alpha}(h,S)]$, our additional unlabeled source data $\mathcal{S}_{3}$ to estimate $\mathbb{E}_{3}[\beta^{2}h^{2}]$, and our unlabeled target data $\mathcal{S}_{2}$ to estimate $\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$. Letting $\hat{\mathbb{E}}_{1}$, $\hat{\mathbb{E}}_{2}$, and $\hat{\mathbb{E}}_{3}$ denote empirical expectations over $\{(X_{i},Y_{i}):i\in[n_{1}]\}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$, respectively, we thus define the empirical risk $\hat{L}(h,\beta)$ as
$\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]+\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h],$ (3)
for $\lambda\geqslant 0$. We then solve the following empirical scaled likelihood ratio regularized quantile regression problem:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}\hat{L}(h,\beta).$ (Empirical-LR-QR)
Our proposed threshold is $q=\hat{h}$. See Algorithm 1.
@@
Input: Labeled source data $(x_{i},y_{i})$, $i\in[n_{1}]$, unlabeled target data $\mathcal{S}_{2}$, unlabeled source data $\mathcal{S}_{3}$, Regularization strength: $\lambda\geqslant 0$, Function class: $\mathcal{H}$
1: Compute $S_{i}=S(x_{i},y_{i})$ for all $i\in[n_{1}]$
2: Solve $(\hat{h},\hat{\beta})\leftarrow$ Equation Empirical-LR-QR
Return: $\hat{C}(x)\leftarrow\{y\in\mathcal{Y}:S(x,y)\leqslant\hat{h}(x)\}$

### v2

@@ Abstract
We consider the problem of conformal prediction under covariate shift. Given labeled data from a source domain and unlabeled data from a covariate shifted target domain, we seek to construct prediction sets with valid marginal coverage in the target domain. Most existing methods require estimating the unknown likelihood ratio function, which can be prohibitive for high-dimensional data such as images. To address this challenge, we introduce the likelihood ratio regularized quantile regression (LR-QR) algorithm, which combines the pinball loss with a novel choice of regularization in order to construct a threshold function without directly estimating the unknown likelihood ratio. We show that the LR-QR method has coverage at the desired level in the target domain, up to a small error term that we can control. Our proofs draw on a novel analysis of coverage via stability bounds from learning theory. Our experiments demonstrate that the LR-QR algorithm outperforms existing methods on high-dimensional prediction tasks, including a regression task for the Communities and Crime dataset, an image classification task from the WILDS repository, and an LLM question-answering task on the MMLU benchmark.
@@ 1 Introduction
  • We demonstrate the effectiveness of the LR-QR algorithm on high-dimensional datasets under covariate shift, including the Communities and Crime dataset, the RxRx1 dataset from the WILDS repository, and the MMLU benchmark. Here, we crucially leverage our theory by choosing the regularization parameter proportional to the theoretically optimal value. An implementation of LR-QR can be accessed at the following link: https://github.com/shayankiyani98/LR-QR.
@@ 1.1 Related work
Here we only list prior work most closely related to our method; we provide more references in Appendix D. The early ideas of conformal prediction were developed in Saunders et al., (1999); Vovk et al., (1999). With the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing prediction sets (e.g., Papadopoulos et al., , 2002; Vovk et al., , 2005; Vovk, , 2013). Classical conformal prediction guarantees validity when the calibration and test data are drawn from the same distribution. In contrast, when there is distribution shift between the calibration and test data (e.g., Quiñonero-Candela et al., , 2009; Shimodaira, , 2000; Sugiyama and Kawanabe, , 2012; Ben-David et al., , 2010; Taori et al., , 2020), coverage may not hold. Covariate shift is a type of dataset shift that arises in many settings, e.g., when predicting disease risk for individuals whose features may evolve over time, while the outcome distribution conditioned on the features remains stable (Quiñonero-Candela et al., , 2009).
@@ 3 Algorithmic principles
Here we present the intuition behind our approach. Our goal is to construct a prediction set of the form $C(x)=\{y\in\mathcal{Y}:S(x,y)\leqslant\tilde{q}(x)\}$, where $\tilde{q}$ should be close to a conditional quantile of $S$ given $X=x$. The quantile loss $\ell_{\alpha}$ is designed such that for any random variable $S$, the minimizers of the objective $\kappa\mapsto\mathbb{E}\ell_{\alpha}(\kappa,S)$ are the $(1-\alpha)$th quantiles of $S$. This has motivated prior work (Jung et al., , 2023; Gibbs et al., , 2025), where the authors minimize the objective $h\mapsto\mathbb{E}\ell_{\alpha}(h(X),S(X,Y))$ for $h$ in some linear hypothesis class $\mathcal{H}$. At a minimizer $h^{*}$, the derivatives in all directions $g\in\mathcal{H}$ should be zero, so that
@@
$\frac{\partial}{\partial\varepsilon}\bigg{|}_{\varepsilon=0}\mathbb{E}_{1}[\ell_{\alpha}(h^{*}(X)+\varepsilon g(X),S)]$ $=\mathbb{E}_{1}[g(X)(\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]=0.$ (2)
If $g$ takes the form $g(x)=d\mathbb{Q}_{X}/d\mathbb{P}_{1,X}(x)$ for some distribution $\mathbb{Q}_{X}$, then [^1] This holds due to the change of measure identity $\mathbb{E}_{P}[dQ/dP(X)h(X)]=\mathbb{E}_{Q}[h(X)]$ for all integrable functions $h$. this equality can be viewed as exact coverage under the covariate shift induced by $g$ for the prediction set $x\mapsto\{y\in\mathcal{Y}:S(x,y)\leqslant h^{*}(x)\}$. In other words, if the test distribution is $\mathbb{Q}=\mathbb{Q}_{X}\times\mathbb{P}_{Y|X}$, then we have the exact coverage result
$\mathbb{E}_{\mathbb{Q}}[\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]=\mathbb{Q}[S(X,Y)\leqslant h^{*}(X)]=1-\alpha.$
Therefore, if the hypothesis class $\mathcal{H}$ is large enough to include the true likelihood ratio $r=d\mathbb{P}_{2,X}/d\mathbb{P}_{1,X}$, then the threshold function $h^{*}$ attains valid coverage in the test domain $\mathbb{P}_{2}$, as desired.
3.1 Our approach
An adaptive choice of the hypothesis class. The above approach requires special assumptions on the hypothesis class $\mathcal{H}$. The choice of the hypothesis class poses a challenge in practice: if $\mathcal{H}$ is too small, then coverage may fail, while if $\mathcal{H}$ is too large, then finite-sample performance may suffer due to large estimation errors.
To address this challenge, our idea is to choose $\mathcal{H}$ adaptively. We start by considering the class of hypotheses $h$ that are close to the true likelihood ratio $r$, as measured by $\mathbb{E}_{1}[(h(X)-r(X))^{2}]$ being small. By our remarks above, if we minimize $\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]$ for $h$ restricted to this set, we obtain a threshold function with valid coverage under the covariate shift $r$.
Removing the explicit dependence on the likelihood ratio. The quantity $\mathbb{E}_{1}[(h(X)-r(X))^{2}]$ depends on the unknown $r$. However, we can expand this to obtain
$\mathbb{E}_{1}[(h(X)-r(X))^{2}]$ $=\mathbb{E}_{1}[h(X)^{2}]+\mathbb{E}_{1}[-2r(X)h(X)]+\mathbb{E}_{1}[r(X)^{2}].$
The term $\mathbb{E}_{1}[r(X)^{2}]$ does not depend on the optimization variable $h$, so it is enough to consider the first two terms. Due to the change-of-measure identity $\mathbb{E}_{1}[r(X)h(X)]=\mathbb{E}_{2}[h(X)]$, the sum of these terms equals
$\mathbb{E}_{1}[h(X)^{2}]+\mathbb{E}_{1}[-2r(X)h(X)]=\mathbb{E}_{1}[h(X)^{2}]+\mathbb{E}_{2}[-2h(X)].$
@@ 3.1 Our approach
A key observation is that neither of the terms $\mathbb{E}_{1}[h(X)^{2}]$ or $\mathbb{E}_{2}[-2h(X)]$ explicitly involve $r$, and thus they can be estimated by sample averages over the source and target data, respectively. Thus, we can minimize $\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]$ over $h\in\mathcal{H}$ while keeping $\mathbb{E}_{1}[h(X)^{2}]+\mathbb{E}_{2}[-2h(X)]$ bounded. The threshold $h^{*}$ will have valid coverage under the covariate shift $r$.
[…]
Replacing the constraint with a regularization. Instead of imposing a constraint on $\min_{\beta\in\mathbb{R}}(\mathbb{E}_{1}[\beta^{2}h(X)^{2}]+\mathbb{E}_{2}[-2\beta h(X)])$, we can use this term as a regularizer. Given a regularization strength $\lambda\geqslant 0$, we can solve
$\min_{h\in\mathcal{H}}\left\{\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]+\lambda\min_{\beta\in\mathbb{R}}(\mathbb{E}_{1}[\beta^{2}h(X)^{2}]+\mathbb{E}_{2}[-2\beta h(X)])\right\}.$
Since the first term does not depend on $\beta$, this is equivalent to the joint optimization problem
$\min_{h\in\mathcal{H},\beta\in\mathbb{R}}\left\{L_{\lambda}(h,\beta):=\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]+\lambda(\mathbb{E}_{1}[\beta^{2}h(X)^{2}]-\mathbb{E}_{2}[2\beta h(X)])\right\}.$ (LR-QR)
@@ 3.2 Algorithm: likelihood ratio regularized quantile regression
We solve an empirical version of this objective. We use our labeled source data $\{(X_{i},Y_{i}):i\in[n_{1}]\}$ to estimate $\mathbb{E}_{1}[\ell_{\alpha}(h(X),S(X,Y))]$, our additional unlabeled source data $\mathcal{S}_{3}$ to estimate $\mathbb{E}_{1}[\beta^{2}h(X)^{2}]$, and our unlabeled target data $\mathcal{S}_{2}$ to estimate $\lambda\mathbb{E}_{2}[-2\beta h(X)]$. Letting $\hat{\mathbb{E}}_{1}$, $\hat{\mathbb{E}}_{2}$, and $\hat{\mathbb{E}}_{3}$ denote empirical expectations over $\{(X_{i},Y_{i}):i\in[n_{1}]\}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$, respectively, we then solve the following empirical likelihood ratio regularized quantile regression problem, for $\lambda\geqslant 0$:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}\left\{\hat{L}_{\lambda}(h,\beta):=\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h(X),S(X,Y))]+\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h(X)^{2}]-\lambda\hat{\mathbb{E}}_{2}[2\beta h(X)]\right\}.$ (Empirical-LR-QR)
Our proposed threshold is $q=\hat{h}$. See Algorithm 1. In the following section, we justify this algorithm through a novel theoretical analysis of the test-time coverage.
@@
Input: $n_{1}$ labeled source datapoints, $n_{2}$ unlabeled target datapoints, $n_{3}$ unlabeled source datapoints
1: Compute scores $S_{i}=S(x_{i},y_{i})$ for all $i\in[n_{1}]$
2: Solve $(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h(X),S(X,Y))]+\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h(X)^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h(X)]$, where $\hat{\mathbb{E}}_{1},\hat{\mathbb{E}}_{2},\hat{\mathbb{E}}_{3}$ denote expectations over the source, unlabeled target, and unlabeled source data;
Return: Prediction set $\hat{C}(x)\leftarrow\{y\in\mathcal{Y}:S(x,y)\leqslant\hat{h}(x)\}$ with asymptotic $1-\alpha$ coverage in the target distribution